Let f(x)=3x^2-9x+7, for what value of x is the value of the cuntion a minimmum? what is the minimum value?
Answer by nerdybill(7384)

Let f(x)=3x^2-9x+7,
for what value of x is the value of the function a minimum?
Since the coefficient associated with the x^2 term is +3, we know that it is a parabola that opens downward. This means the "axis of symmetry" will be the minimum.
x = -b/2a = -(-9)/(2*3) = 9/6 = 3/2 = 1.5
.
what is the minimum value?
f(x)=3x^2-9x+7
f(1.5)=3(1.5^2)-9(1.5)+7
f(x)= .25